# A Survey on LLM-as-a-Judge

### v1

@@ 8.3. MLLM-as-a-Judge
AI systems are rapidly evolving into highly multifunctional entities. For example, whereas in the past we had special-purpose solutions for different language processing tasks (e.g., sentiment analysis, parsing, dialogue), LLMs are competent at all these tasks using a single set of weights (Srivastava et al., 2022). Unified systems are also being built across data modalities: instead of using a different architecture for processing images versus text, recent models, such as GPT4-V (OpenAI, 2023b), Gemini (Google, 2023), and LLaVA (Liu et al., 2023b), handle both modalities In brief, AI systems are becoming more uniform in their structures and functions. This also applies to LLM-as-a-Judge.
The future of evaluation lies in developing robust multi-modal evaluators that can handle complex, multi-modal content, such as text, images, and video. Current multi-modal LLMs, while promising, lack the reasoning depth and reliability of their text-based counterparts. Future research should focus on:
A practical multi-modal evaluator will not only advance research but also open new possibilities for tasks such as multi-modal content moderation and knowledge extraction.

### v2

@@ 8.3. MLLM-as-a-Judge
AI systems are evolving into highly versatile and multifunctional entities. Traditionally, specialized models were required for distinct language processing tasks, such as sentiment analysis, syntactic parsing, and dialogue modeling. However, large language models (LLMs) have demonstrated competence across these tasks using a single set of weights (Srivastava et al., 2022). Similarly, advancements are being made toward unified systems capable of processing multiple data modalities. Instead of employing distinct architectures for processing text, audio, and images, recent models like GPT-4o (OpenAI, 2023b), Gemini (Google, 2023), and LLaVA (Liu et al., 2023c) integrate these capabilities within a single framework. These developments highlight a growing trend toward unification in the structure and functionality of AI systems, which extends to the emerging paradigm of LLM-as-a-Judge.
Currently, the emergence of MLLM-as-a-Judge frameworks (Chen et al., 2024c) for evaluating models can be observed. However, research exploring how MLLM-as-a-Judge could be applied to the evaluation of data or agents remains limited. Beyond model evaluation, MLLM-as-a-Judge, much like LLM-as-a-Judge, is envisioned to have the capability to assess or annotate data, function as a Reward Model, or serve as a Verifier within intermediate reasoning processes. These expanded roles would allow MLLM-as-a-Judge to contribute more broadly to the AI pipeline.
The future of evaluation lies in developing robust multimodal evaluators capable of reasoning about and assessing complex content spanning text, audio, images, and video. While current multimodal LLMs exhibit promising capabilities, they often lack the reasoning depth and reliability of their text-based counterparts. Future research must address these limitations, with a focus on enhancing reasoning capabilities, improving reliability, and enabling seamless integration across modalities. A practical multimodal evaluator has the potential to not only advance AI research but also enable new applications in areas such as multimodal content moderation and automated knowledge extraction.
@@ 8.4. More LLM-as-a-Judge Benchmarks
8.5. LLM-as-a-Judge for LLM Optimization
LLM-as-a-Judge shows substantial promise for advancing LLM optimization. Recent studies (Zhuge et al., 2024) have begun incorporating LLM-as-a-Judge into multi-agent frameworks to guide inter-agent interactions, thereby improving overall decision-making efficiency and quality. In addition, LLM-as-a-Judge has been employed in Reinforced Fine-Tuning (ReFT) pipelines (Trung et al., 2024), functioning as a crucial scoring module for evaluating the reasoning processes of models. By flexibly adapting to diverse content formats and domains, LLM-as-a-Judge offers a robust and efficient evaluation mechanism for a wide range of optimization tasks.
Despite these encouraging developments, current research efforts are still in their infancy. Future work should focus on broadening the application domains and strategies for implementing LLM-as-a-Judge, especially in complex, multimodal scenarios. Furthermore, a systematic assessment of its reliability and generalization capabilities will be critical for fully realizing the potential of LLM-as-a-Judge in enhancing model performance and robustness.
9. Conclusion
LLM-as-a-Judge represents an LLM-based paradigm for evaluation, offering a compelling alternative to traditional methods that rely either on human experts or quantitative assessments. By leveraging the powerful capabilities of LLMs, this framework excels in scoring, rating, and evaluating tasks with exceptional scalability, adaptability, and precision. These qualities align directly with the growing demand for efficient and robust evaluation systems across diverse fields.
However, fully realizing the potential of LLM-as-a-Judge requires addressing significant challenges, particularly those concerning reliability. The evaluation process must account for factors such as consistency, bias mitigation, and contextual adaptability, ensuring that the system operates under well-defined rules while delivering human-aligned and objective results. Each stage of development—from dataset curation and model fine-tuning to evaluation protocol standardization—demands careful consideration to uphold reliability. Simply deploying LLM-as-a-Judge does not guarantee outcomes that are both accurate and aligned with evaluation needs. Building reliable systems requires systematic efforts to evaluate and refine its outputs, while addressing gaps in definitions, practices, and research.
This survey has provided a comprehensive review of the LLM-as-a-Judge landscape, offering insights into its current state and practical strategies for improvement. By addressing these challenges through robust methodologies and innovative benchmarks, LLM-as-a-Judge can evolve into a highly reliable and practical tool for diverse applications. Ultimately, while the current capabilities of LLMs are impressive, their effectiveness as evaluators depends on thoughtful design and systematic refinement. We hope this work serves as a foundation for further research, guiding the development of more reliable LLM-as-a-Judge systems and facilitating their widespread deployment across academic, industrial, and societal domains.